OpenStudy (anonymous):

Which equation would best help solve the following problem? Maria releases a javelin 1.5 meters above the ground with an initial vertical velocity of 22 meters per second. How long will it take the javelin to hit the ground?

OpenStudy (anonymous):

a –4.9t^2 + 22t + 1.5 = 0 b –4.9t^2+ 22t = 1.5 c –4.9t^2 + 1.5t + 22 = 0 d –16t^2 + 22t + 1.5 = 0